A small company has an advertising budget of $15,000. The company plans to produce and air a television commercial. It will cost $500 to produce the commercial and an additional $50 each time the commercial is aired. How many days can the company afford to run the commercial if it aired once a day?
15,000- 500 14500 divided by 50 290
Answer by checkley77(12569)

(15,000-500)/50=14,500/50=290 days can be payed for.